# Extending Stable and Popular Matching Algorithms from Bipartite to Arbitrary Instances

### v1

@@ 3.2 Maximum Size Popular Fractional-Matching
Let this cycle be $C^{k}$. Also observe that by the construction of $H^{k}$, the edges $C^{k}\cap F^{k}$ alternate between $(e_{M}^{i},e_{M}^{j})$ type and $(e_{N}^{i},e_{N}^{j})$ type edges. Hence, $C^{k}\cap F^{k}$ gives a walk $P^{k}$ in $G$, where $P^{k}=(e_{0},f_{1},e_{1},\dots,f_{j_{k}},e_{j_{k}})$, with $M^{k}(e_{i})>0$, $N^{k}(f_{i})>0$ and $\phi_{v_{i-1}}(e_{i-1},f_{i})>0,\phi_{u_{i}}(e_{i},f_{i})>0$ and it may happen that $e_{0}=\emptyset$, $e_{j_{k}}=\emptyset$ or $e_{0}=e_{j_{k}}$ (note that an edge can have $N^{k}(e)>0$ and $M^{k}(e)>0$ simultaneously). Here, $e_{i}=(u_{i},v_{i})$ and $f_{i}=(v_{i-1},u_{i})$.
Set the weight of $P_{k}$ to be
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i}),\phi_{u_{i}}^{k}(e_{i},f_{i})\}$, if $e_{0}=e_{j_{k}}$,
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i}),\phi_{u_{i}}^{k}(e_{i},f_{i}),\phi_{u_{0}}(e_{0},\emptyset)\}$ if $e_{0}\neq\emptyset,e_{j_{k}}=\emptyset$,
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i}),\phi_{u_{i}}^{k}(e_{i},f_{i}),\phi_{v_{j_{k}}}(e_{j_{k}},\emptyset)\}$, if $e_{0}=\emptyset,e_{j_{k}}\neq\emptyset$ and finally
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i}),\phi_{u_{i}}^{k}(e_{i},f_{i}),\phi_{u_{0}}(e_{0},\emptyset),\phi_{v_{j_{k}}}(e_{j_{k}},\emptyset)\}$, if $\emptyset\neq e_{0}\neq e_{j_{k}}\neq\emptyset$.
Then, we let $k:=k+1$, $\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})=\phi_{v_{i-1}}^{k-1}(e_{i-1},f_{i})-w_{k-1}$, $\phi^{k}_{u_{i}}(e_{i},f_{i})=\phi_{u_{i}}^{k-1}(e_{i},f_{i})-w_{k-1}$ and $\phi_{v}^{k}(e,f)=\phi_{v}^{k-1}(e,f)$ otherwise. If $v^{\emptyset}\in C^{k-1}$, then we let $\phi^{k}_{v_{i}}(e,\emptyset)=\phi^{k-1}_{v_{i}}(e,\emptyset)-w_{k-1}$ if $(v^{\emptyset},e^{i})\in C^{k-1}$ and $e=e_{0},v_{i}=u_{0}$ or $e=e_{j_{k}},v_{i}=v_{j_{k}}$, and $\phi^{k}_{v}(e,\emptyset)=\phi^{k-1}_{v}(e,\emptyset)$ otherwise. Also, $\phi_{v}^{k}(\emptyset,f)=\phi^{k-1}_{v}(\emptyset,f)-w_{k-1}$, if $e_{0}=\emptyset,f=f_{1},v=v_{0}$ or $e_{j_{k}}=\emptyset,f=f_{j_{k}},v=u_{j_{k}}$ and $\phi^{k}_{v}(\emptyset,e)=\phi^{k-1}_{v}(\emptyset,e)$ otherwise.
We let $M^{k}(e_{i})=M^{k-1}(e_{i})-w_{k-1}$ and $M^{k}(e)=M^{k-1}(e)$ otherwise. We let $N^{k}(f_{i})=N^{k-1}(f_{i})-w_{k-1}$ and $N^{k}(f)=N^{k-1}(f)$ otherwise. This is feasible, as each edge appears at most once in as an $e_{i}$ edge and at most once as an $f_{i}$ edge in $P^{k}$.
@@
Finally, we define $H^{k}$ from $M^{k},N^{k}$ and $\phi^{k}$ the same way and iterate this procedure until $\phi^{k}\equiv 0$. We can iterate it, since each vertex we keep in $H^{k}$ apart from $v^{\emptyset}$ remains adjacent to at least one edge from both $F^{k}$ and $E^{k}\setminus F^{k}$ as we maintain $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}^{k}(e,f)=\sum_{f\in E(u)\cup\{\emptyset\}}\phi_{u}^{k}(e,f)=M^{k}(e)$ and $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}^{k}(f,e)=\sum_{f\in E(u)\cup\{\emptyset\}}\phi_{u}^{k}(f,e)=N^{k}(e)$ (in both cases, either all three terms remain the same, or all three decreases by $w_{k-1}$ and by the choice of $w_{k-1}$ it remains nonnegative). This procedure terminates in finitely many steps, as in each step at least one new $\phi^{k}_{v}(e,f)$ becomes 00.
@@
For each walk $P^{l}$, let $P^{l}=(e_{0},f_{1},e_{1},f_{2},e_{2}\dots,e_{j_{l}})$, where $M(e_{i})>0$ and $N(f_{i})>0$ for each $i$ and $e_{0}=e_{j_{l}}$, if $P^{l}$ is a closed walk and $e_{0},e_{j_{l}}$ can be $\emptyset$.
@@
Let $e_{i}=(u_{i},v_{i})$ and $f_{i}=(v_{i-1},u_{i})$. Of course, it may happen that $v_{i}=v_{j}$ or $u_{i}=v_{j}$ for some $i,j$, however, this does not matter for the proof. Note that depending on the above cases, we can have $u_{0}=\emptyset$, $v_{j_{l}}=\emptyset$ or $u_{0}=u_{j_{l}},v_{0}=v_{j_{l}}$.
@@
Then, we have that $\Delta(M,N,\phi)=\sum\limits_{l=1}^{k}w_{k}\cdot\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))$, where $x(P^{l})=\mathrm{vote}_{u_{0}}(e_{0},\emptyset)+\mathrm{vote}_{v_{j_{l}}}(e_{j_{l}},\emptyset)$ if $e_{0}\neq e_{j_{l}}$ and 0 otherwise. Therefore, it is enough to show that for each walk $P^{l}$, $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$ and it is strictly larger than 00, whenever we have $e_{0}=e_{j_{l}}=\emptyset$.
@@
    If $e_{i}$ has a bad copy for $v_{i}$ with positive value in $M^{\prime}$ and $e_{i+1}$ has a good copy for $v_{i+1}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})=+2$.
@@
    If $e_{i}$ has a bad copy for $v_{i}$ with positive value in $M^{\prime}$ or $e_{i+1}$ has a good copy for $v_{i+1}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})\geq 0$.
@@
    If $e_{i}$ has a good copy for $v_{i}$ with positive value in $M^{\prime}$ and $e_{i+1}$ has a bad copy for $v_{i+1}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})\geq-2$.
@@
2. In this case, at least one of $u_{i+1},v_{i}$ is matched with a bad copy with positive weight in $M^{\prime}$, say it is $v_{i}$ (other case is analogous). Hence, if both prefers $f_{i+1}$, then the bad copy of $f_{i+1}$ for $v_{i}$ (unsaturated in $M^{\prime}$ for the same reasons) blocks $M^{\prime}$, contradiction.
@@
4. If $e_{0}=\emptyset$, then we must have by the 3-rd axiom of sensible pairings that $v_{0}$ is unsaturated in $M^{\prime}$, as $\phi_{v_{0}}(\emptyset,f_{1})>0$. Hence, by the fact that the bad copy of $f_{1}$ for $v_{0}$ does not block $M^{\prime}$, we get that $u_{1}$ is saturated with good copies in $M^{\prime}$, so $e_{1}$ must have a bad copy with positive weight in $M^{\prime}$ for $v_{1}$. Also, $u_{1}$ has to prefer each edge of $M^{\prime}$ with positive weight even to the good copy of $f_{1}$, so $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
5. This case is very similar to 3. We have that $u_{j_{l}}$ is unsaturated, so $v_{j_{l}-1}$ can only have good copies with positive weight in $M^{\prime}$ and it must prefer $e_{j_{l}-1}$ to $f_{j_{l}}$, so $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$. $\blacksquare$
Using $\mathrm{vote}_{v_{i}}(e_{i},f_{i})+\mathrm{vote}_{u_{i+1}}(e_{i},f_{i+1})\geq-2$, we get by the above observations that for each $P^{l}$, case 1. happens at least as many times as case 3., except maybe in the case, where $e_{0}\neq\emptyset$, $e_{j_{l}}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$, in which case it can happen once more. However, in that case, the two endpoints $u_{0}$ and $v_{j_{l}}$ vote with $+1$, i.e. $x(P^{l})=+2$, so $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$.
Finally, if $e_{0}=e_{j_{l}}=\emptyset$, then case 1. happens one more time than case 3., so by cases 4. and 5. we have that $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))>0$. This proves that $M$ is extra-popular and that it has a maximum size even among the barely-defendable fractional-matchings. ∎
@@
Observe that we did not use the 4-th axiom of sensible pairings anywhere in the proof of Theorem 3. Hence, the half-matching we get is extra-popular even in a slightly stronger sense we get by eliminating axiom 4 from the definition of sensible pairings. Also, it is largest among barely-defendable fractional-matchings even with the weaker definition which does not require axiom 4 from sensible pairings.
The reason for axiom 4 is that it is quite natural and symmetric to axiom 3, and that we utilize it in the next section for the Pop-Crit-SRI algorithm.

### v3

@@ 3.2 Maximum Size Popular Fractional-Matching
Let this closed walk be $C^{k}$. Also observe that by the construction of $H^{k}$, the edges $C^{k}\cap F^{k}$ alternate between $(e_{M}^{i},e_{M}^{j})$ type and $(e_{N}^{i},e_{N}^{j})$ type edges. Hence, $C^{k}\cap F^{k}$ gives a walk $P^{k}$ in $G$, where $P^{k}=(e_{0},f_{1},e_{1},\dots,f_{j_{k}},e_{j_{k}})$, with $M^{k}(e_{i})>0$, $N^{k}(f_{i})>0$ and $\phi_{v_{i-1}}(e_{i-1},f_{i})>0,\phi_{u_{i}}(e_{i},f_{i})>0$ and it may happen that $e_{0}=\emptyset$, $e_{j_{k}}=\emptyset$ or $e_{0}=e_{j_{k}}$ (note that an edge can have $N^{k}(e)>0$ and $M^{k}(e)>0$ simultaneously).
Intuitively, what we do in the following paragraphs is decreasing the $M^{k},N^{k}$ and $\phi^{k}$ values according to the $e_{i},f_{i}$ edges and $(e_{i-1},f_{i}),(e_{i},f_{i})$ comparisons that are given by the closed walk $C^{k}$, until one new $\phi^{k}(e,f)$ becomes 0 (while also being careful about the possibly different multiplicities of these in $C^{k}$).
Let $e_{i}=(u_{i},v_{i})$ and $f_{i}=(v_{i-1},u_{i})$. Let $\alpha_{i-1,i}$ be the multiplicity of the edge of $E^{k}$ corresponding to the comparison $(e_{i-1},f_{i})$ in $C^{k}$ (the edge corresponding to a comparison $(e,f)$ is $(e_{M}^{i},f_{N}^{i})$), $\alpha_{i,i}$ be the multiplicity of the edge of $E^{k}$ corresponding to the comparison $(e_{i},f_{i})$ in $C^{k}$, $\alpha_{i}$ be the multiplicity of the edge of $F^{k}$ corresponding to the $M$-version of $e_{i}$ (the $M$-version of an edge $e$ is $(e_{M}^{i},e_{M}^{j})$) in $C^{k}$ and finally $\beta_{i}$ be the multiplicity of the edge of $F^{k}$ corresponding to the $N$-version of $f_{i}$ in $C^{k}$. Similarly, let $\alpha_{0,0}$ be the multiplicity of the edge of $E^{k}$ corresponding to the comparison $(e_{0},\emptyset)$ in $C^{k}$, if $v^{\emptyset}\in C^{k}$, but $e_{0}\neq\emptyset$ and $\alpha_{j_{k},0}$ be the multiplicity of the edge of $E^{k}$ corresponding to the comparison $(e_{j_{k}},\emptyset)$ in $C^{k}$, if $v^{\emptyset}\in C^{k}$, but $e_{j_{k}}\neq\emptyset$.
Set the weight of $P^{k}$ to be
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})/\alpha_{i-1,i},\phi_{u_{i}}^{k}(e_{i},f_{i})/\alpha_{i,i}\}$, if $e_{0}=e_{j_{k}}$,
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})/\alpha_{i-1,i},\phi_{u_{i}}^{k}(e_{i},f_{i})/\alpha_{i,i},\phi_{u_{0}}(e_{0},\emptyset)/\alpha_{0,0}\}$ if $e_{0}\neq\emptyset,e_{j_{k}}=\emptyset$,
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})/\alpha_{i-1,i},\phi_{u_{i}}^{k}(e_{i},f_{i})/\alpha_{i,i},\phi_{v_{j_{k}}}(e_{j_{k}},\emptyset)/\alpha_{j_{k},0}\}$, if $e_{0}=\emptyset,e_{j_{k}}\neq\emptyset$ and finally
  • $w_{k}=\min_{i\in[j_{k}]}\{\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})/\alpha_{i-1,i},\phi_{u_{i}}^{k}(e_{i},f_{i})\alpha_{i,i},\phi_{u_{0}}(e_{0},\emptyset)/\alpha_{0,0},\phi_{v_{j_{k}}}(e_{j_{k}},\emptyset)/\alpha_{j_{k},0}\}$, if $\emptyset\neq e_{0}\neq e_{j_{k}}\neq\emptyset$.
Then, we let $k:=k+1$, $\phi_{v_{i-1}}^{k}(e_{i-1},f_{i})=\phi_{v_{i-1}}^{k-1}(e_{i-1},f_{i})-\alpha_{i-1,i}w_{k-1}$, $\phi^{k}_{u_{i}}(e_{i},f_{i})=\phi_{u_{i}}^{k-1}(e_{i},f_{i})-\alpha_{i,i}w_{k-1}$ ($i=1,\dots,j_{k-1}$) and $\phi_{v}^{k}(e,f)=\phi_{v}^{k-1}(e,f)$ otherwise. If $v^{\emptyset}\in C^{k-1}$ but $e_{0}\neq\emptyset$, then we let $\phi^{k}_{u_{0}}(e_{0},\emptyset)=\phi^{k-1}_{u_{0}}(e_{0},\emptyset)-\alpha_{0,0}w_{k-1}$, $\phi^{k}_{v_{j_{k-1}}}(e_{j_{k}},\emptyset)=\phi_{v_{j_{k-1}}}^{k-1}(e_{j_{k}},\emptyset)-\alpha_{j_{k-1},0}w_{k-1}$ and $\phi^{k}_{v}(e,\emptyset)=\phi^{k-1}_{v}(e,\emptyset)$ otherwise.
We let $M^{k}(e_{i})=M^{k-1}(e_{i})-\alpha_{i}w_{k-1}$ ($i=0,1,\dots,j_{k-1}$) and $M^{k}(e)=M^{k-1}(e)$ otherwise. We let $N^{k}(f_{i})=N^{k-1}(f_{i})-\beta_{i}w_{k-1}$ ($i=1,\dots,j_{k-1}$) and $N^{k}(f)=N^{k-1}(f)$ otherwise. By the choice of $w_{k-1}$, every value remains nonnegative.
@@
Finally, we define $H^{k}$ from $M^{k},N^{k}$ and $\phi^{k}$ the same way and iterate this procedure until $\phi^{k}\equiv 0$. We can iterate it, since each vertex we keep in $H^{k}$ apart from $v^{\emptyset}$ remains incident to at least one edge from both $F^{k}$ and $E^{k}\setminus F^{k}$ as the value updates maintain $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}^{k}(e,f)=\sum_{f\in E(u)\cup\{\emptyset\}}\phi_{u}^{k}(e,f)=M^{k}(e)$ and $\sum_{f\in E(v)\cup\{\emptyset\}}\phi_{v}^{k}(f,e)=\sum_{f\in E(u)\cup\{\emptyset\}}\phi_{u}^{k}(f,e)=N^{k}(e)$. This procedure terminates in finitely many steps, as in each step at least one new $\phi^{k}_{v}(e,f)$ value becomes 0.
@@
For each walk $P^{l}$, let $P^{l}=(e_{0},f_{1},e_{1},f_{2},e_{2}\dots,e_{j_{l}})$, where $M(e_{i})>0$ and $N(f_{i})>0$ for each $i$ and $e_{0}=e_{j_{l}}$, if $P^{l}$ is a closed walk. Note that $e_{0},e_{j_{l}}$ can be $\emptyset$.
@@
Recall that $e_{i}=(u_{i},v_{i})$ and $f_{i}=(v_{i-1},u_{i})$. Of course, it may happen that $v_{i}=v_{j}$ or $u_{i}=v_{j}$ for some $i,j$, however, this does not matter for the proof. Note that depending on the above cases, we can have $u_{0}=\emptyset$, $v_{j_{l}}=\emptyset$ or $u_{0}=u_{j_{l}},v_{0}=v_{j_{l}}$.
@@
Then, we have that $\Delta(M,N,\phi)=\sum\limits_{l=1}^{k}w_{l}\cdot\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))$, where $x(P^{l})=\mathrm{vote}_{u_{0}}(e_{0},\emptyset)+\mathrm{vote}_{v_{j_{l}}}(e_{j_{l}},\emptyset)$ if $e_{0}\neq e_{j_{l}}$ and 0 otherwise. Therefore, it is enough to show that for each walk $P^{l}$, $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$ and it is strictly larger than 0, whenever we have $e_{0}=e_{j_{l}}=\emptyset$.
@@
    If $e_{i}$ has a bad copy for $v_{i}$ with positive value in $M^{\prime}$ and $e_{i+1}$ has a good copy for $v_{i+1}$ with positive value in $M^{\prime}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})=+2$.
@@
    If $e_{i}$ has a bad copy for $v_{i}$ with positive value in $M^{\prime}$ or $e_{i+1}$ has a good copy for $v_{i+1}$ with positive value in $M^{\prime}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})\geq 0$.
@@
    If $e_{i}$ has a good copy for $v_{i}$ with positive value in $M^{\prime}$ and $e_{i+1}$ has a bad copy for $v_{i+1}$ with positive value in $M^{\prime}$, then $\mathrm{vote}_{v_{i}}(e_{i},f_{i+1})+\mathrm{vote}_{u_{i+1}}(e_{i+1},f_{i+1})\geq-2$.
@@
2. In this case, at least one of $u_{i+1},v_{i}$ is matched with a bad copy with positive weight in $M^{\prime}$, say it is $v_{i}$ (the other case is analogous). Hence, if both prefers $f_{i+1}$, then the bad copy of $f_{i+1}$ for $v_{i}$ (unsaturated in $M^{\prime}$ for the same reasons) blocks $M^{\prime}$, contradiction.
@@
4. If $e_{0}=\emptyset$, then we must have by axiom 3 of sensible pairings that $v_{0}$ is unsaturated in $M^{\prime}$, as $\phi_{v_{0}}(\emptyset,f_{1})>0$. Hence, by the fact that the bad copy of $f_{1}$ for $v_{0}$ does not block $M^{\prime}$, we get that $u_{1}$ is saturated with good copies in $M^{\prime}$, so $e_{1}$ must have a bad copy with positive weight in $M^{\prime}$ for $v_{1}$. Also, $u_{1}$ has to prefer each edge of $M^{\prime}$ with positive weight even to the good copy of $f_{1}$, so $\mathrm{vote}_{u_{1}}(e_{1},f_{1})=+1$.
5. This case is very similar to 4. We have that $u_{j_{l}}$ is unsaturated, so $v_{j_{l}-1}$ can only have good copies with positive weight in $M^{\prime}$ and it must prefer $e_{j_{l}-1}$ to $f_{j_{l}}$, so $\mathrm{vote}_{v_{j_{l}-1}}(e_{j_{l}-1},f_{j_{l}})=+1$.  $\blacksquare$
We get by the above observations that for each $P^{l}$, case 1 happens at least as many times as case 3, except when $e_{0}\neq\emptyset$, $e_{j_{l}}\neq\emptyset$ and $e_{0}\neq e_{j_{l}}$, in which case it can happen once more. However, then the two endpoints $u_{0}$ and $v_{j_{l}}$ vote with +1+1, i.e. $x(P^{l})=+2$, so $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))\geq 0$ holds regardless.
Finally, if $e_{0}=e_{j_{l}}=\emptyset$, then case 1 happens one more time than case 3 by 4 and 5, so we have that $\sum\limits_{j=1}^{j_{l}}(\mathrm{vote}_{v_{j-1}}(e_{j-1},f_{j})+\mathrm{vote}_{u_{j}}(e_{j},f_{j})+x(P^{l}))>0$. This proves that $M$ is extra-popular and that it has a maximum size even among the barely-defendable fractional-matchings. ∎
@@
Observe that we did not use axiom 4 of sensible pairings anywhere in the proof of Theorem 3. Hence, the half-matching we get is extra-popular even in a slightly stronger sense we get by eliminating axiom 4 from the definition of sensible pairings. Also, it is largest among barely-defendable fractional-matchings even with the weaker definition which does not require axiom 4 from sensible pairings.
The reason for axiom 4 is that it is quite natural and symmetric to axiom 3, and that we utilize it in the next section for the Crit-fPRI algorithm.